Adriana took 2 h longer to drive 360 mi on the first day of a trip than she took to drive 270 mi on the second day. If her speed was the same on both days, what was the driving time each day?
On this one I came up with the first day 4 hours and 12 min and the second day I came up with 3 hours and 20 min. Is this correct. Thanks

Answer by checkley75(3666)

360/R=270/R+2
360/R=(270+2R)/R NOW CANCEL OUT THE DENOMINATORS & WE HAVE LEFT
360=270+2R
2R=360-270
2R=90
R=90/2
R=45 MPH FOR THE RATE FOR BOTH TRIPS.
PROOF
360/45=270/45+2
8=6+2
8=8